A firm has a market capitalization (market value of equity) of $19 Billion and net debt of $11 Billion. Calculate the weight of equity in the firm's weighted average cost of capital (WACC) calculation. [Note: Enter your answer as a percentage rounded to two decimal places.]
Answer #1

Weight of equity = $19 Billion / ($19 Billion + $11 Billion)

Weight of equity = $19 Billion / $30 Billion

Weight of equity = 0.6333 or 63.33%
Answer #2

ANSWER :




C = E + D = 19 + 11 = 30 billion dollars.


Weight of equity in the capital = E/C = 19 / 30 = 0.6333 = 63.33% (ANSWER).

answered by: Tulsiram Garg